If in a stick of chalk, carbon is 3g, what is the weight of the chalk stick?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Weight of carbon in the chalk = 3g
Let us assume the weight of the stick be x
Then,
12% of x = 3
(12/100) × (x) = 3
X = 3 × (100/12)
X = 1 × (100/4)
X = 25g
∴The weight of the stick is 25g.

Thanks